Show that xp-px=ih for the ground state wave function of the quantum harmonic oscillator.


1
Expert's answer
2021-02-25T17:28:15-0500

Consider "\hat A = \hat x" and Bˆ = "\hat p = −ih\dfrac{d}{dx}" ,


Then we have "\hat A\hat Bf(x) = \hat x\hat pf (x)"


We can of course also construct another new operator : "\hat p\hat x"


Then, by definition of the operator product, "\hat p\hat xf (x)"


means that "\hat x" is first operating on f(x) and then "\hat p" is operating on the function "\hat xf (x)" .


Compare the results of operating with the products "\hat p\hat x \text{ and } \hat x\hat \ p \text{ on } f(x)"


"(\hat x\hat p− \hat p\hat x)f(x) = −ih(x\dfrac{d(f(x))}{dx} - \dfrac{d}{dx}(xf(x)))"  


and hence by the product rule of differentiation:


"(\hat x\hat p− \hat p\hat x)f(x) = ihf(x)"


and since this must hold for any differentiable function f(x), we can write this as an operator equation:


"\hat x\hat p− \hat p\hat x = ih"


Thus, we have shown that the operator product of "\hat x" and "\hat p" is non-commuting.